# Learning Where to Learn: Training Distribution Selection for Provable OOD Performance

### v1

@@ E.1 Bilevel minimization for function approximation
$\mathbb{Q}\coloneqq\frac{1}{K}\sum_{k=1}^{K}\delta_{\nu_{k}^{\prime}}\,.$ (E.6)
@@
The test measures $\nu_{k}^{\prime}\in\mathscr{P}_{2}(\mathbb{R}^{d})$ in (E.6) are deterministic and fixed. In applications, these could be the datasets for which we require good OOD accuracy. However, to synthetically generate the $\{\nu_{k}^{\prime}\}$ in practice, we set
@@
$\nu_{k}^{\prime}\coloneqq\mathcal{N}(m_{k}^{\prime},\mathcal{C}_{k}^{\prime})\,,\quad\text{where}\quad m_{k}^{\prime}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,I_{d})\quad\text{and}\quad\mathcal{C}^{\prime}_{k}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathsf{Wishart}_{d}(I_{d},d+1)$ (E.7)
@@
for each $k=1,\ldots,K$. Once the $K$ realizations of these distributions are generated, we view them as fixed once and for all. We assume that $K$ and $\mathbb{Q}$ are fixed, but we only have access to the empirical measure $\widehat{\mathbb{Q}}$ defined by (E.6) but with each $\nu_{k}^{\prime}$ replaced by the empirical measure $\frac{1}{M}\sum_{m=1}^{M}\delta_{v_{m}^{(k)}}$ generated by samples $\{v_{m}^{(k)}\}_{m=1}^{M}\sim{\nu_{k}^{\prime}}^{\otimes M}$. In the experiments, we set $K=10$ and $M=5000$. To create a fair test set, we use a fixed $500$ sample validation subset to update the adjoint state as in Lem. C.5 and to monitor the validation OOD accuracy during gradient descent. The remaining $4500$ samples are used to evaluate the final OOD test accuracy of our trained models; see (E.8).
@@
It remains to describe the optimization of the Gaussian family $\nu_{\vartheta}=\mathcal{N}(m_{0},\mathcal{C})$. Practically, we actually optimize over the lower triangular Cholesky factor $\vartheta\coloneqq L$ of $\mathcal{C}=LL^{\top}\in\mathbb{R}^{d\times d}$ instead of $\mathcal{C}$ itself. This removes redundant parameters due to the symmetry of $\mathcal{C}$. We enforce strict positive definiteness of $\mathcal{C}$ by replacing all nonpositive diagonal entries of $L$ with a threshold value of $10^{-7}$. Other than this constraint, each entry of $L$ is optimized over the whole of $\mathbb{R}$. Recalling the log density formula (4.5) from Ex. 4.5, we use automatic differentiation to compute $\nabla_{\vartheta}\log(p_{\vartheta})$; this factor appears in the definition of the approximate gradient (C.6) and in Alg. 3.
@@
$\mathsf{Err}\coloneqq\sqrt{\frac{\operatorname{\mathbb{E}}_{\nu^{\prime}\sim\widehat{\mathbb{Q}}}\operatorname{\mathbb{E}}_{u^{\prime}\sim\nu^{\prime}}\big{\lvert}\mathcal{G}^{\star}(u^{\prime})-\widehat{\mathcal{G}}_{N}^{(\nu,\sigma)}(u^{\prime})\big{\rvert}^{2}}{\operatorname{\mathbb{E}}_{\nu^{\prime}\sim\widehat{\mathbb{Q}}}\operatorname{\mathbb{E}}_{u^{\prime}\sim\nu^{\prime}}\big{\lvert}\mathcal{G}^{\star}(u^{\prime})\big{\rvert}^{2}}}\approx\sqrt{\frac{\mathscr{E}_{\mathbb{Q}}(\widehat{\mathcal{G}}_{N}^{(\nu,\sigma)})}{\mathscr{E}_{\mathbb{Q}}(0)}}\,.$ (E.8)
@@
In the preceding display, we observe that $\mathsf{Err}$ employs a Monte Carlo discretization of the integrals appearing in $\mathscr{E}_{\mathbb{Q}}$ by averaging over the empirical meta measure $\widehat{\mathbb{Q}}$, which is composed of $K=10$ empirical test measures. Each individual test measure itself is an empirical measure with $4500$ atoms at points that are unseen during the adjoint state calculation. See the preceding paragraph for details about our choice of $\mathbb{Q}$. We report values of $\mathsf{Err}$ in Figs. 2 and 7 and in Tab. 1.
@@
E.2 AMINO for learning electrical impedance tomography
@@
To illustrate this, we consider the EIT problem with the AMINO solver; see SM D.3. Suppose the general distribution of interest for the Dirichlet boundary condition is generated according to the law of $f_{l}$ in (D.4) with $\omega\sim\mathsf{Unif}([\frac{\pi}{2},7\pi])$. A reasonable distribution to train upon is this very same distribution. However, suppose we were constrained to focus and train on distributions corresponding to $\operatorname{Law}(\omega)$ supported only on a subset of the domain $[\frac{\pi}{2},7\pi]$. In that case, the choice of the distribution will greatly affect the OOD generalization error when testing with $\omega\sim\mathsf{Unif}([\frac{\pi}{2},7\pi])$. To be more concrete, we will focus on four training distributions for $\omega$, which directly determines the input Dirichlet boundary condition as in SM D.3: $\mathsf{Unif}([\frac{\pi}{2},\pi])$, $\mathsf{Unif}([\frac{7\pi}{2},4\pi])$, $\mathsf{Unif}([5\pi,\frac{11\pi}{2}])$, and $\mathsf{Unif}([\frac{13\pi}{2},7\pi])$. After obtaining the well-trained neural operators (in the sense that the ID generalization error is small enough) corresponding to each training distribution, we study the OOD generalization error of the trained neural operators for Dirichlet data obtained by drawing $\omega\sim\mathsf{Unif}([\frac{\pi}{2},7\pi])$ in (D.4).
@@
We train four distinct AMINO models; see SM D.3 to recall AMINO and SM D.1 for its DeepONet component. Each model comprises three jointly trained components:
@@
  • Trunk network: An $8$–layer multilayer perceptron (MLP) with $512$ neurons per hidden layer, LeakyReLU activations, zero dropout, and a final output dimension of $p=100$.
@@
The four neural operator models “Model 1”, “Model 2”, “Model 3”, and “Model 4” are trained using four different input parameter distributions for $\operatorname{Law}(\omega)$, respectively: $\mathsf{Unif}([\frac{\pi}{2},\pi])$, $\mathsf{Unif}([\frac{7\pi}{2},4\pi])$, $\mathsf{Unif}([5\pi,\frac{11\pi}{2}])$, and $\mathsf{Unif}([\frac{13\pi}{2},7\pi])$; see (D.4) to recall how $\omega$ is reflected in the actual input distributions on function space. For each distribution, $4096$ training samples are generated. Each sample consists of $\mathsf{L}=100$ Dirichlet–Neumann boundary condition pairs with the underlying conductivity discretized on a $70\times 70$ grid for the domain $\Omega=(0,1)^{2}$. After $500$ training epochs, the trained models are evaluated on $1024$ test samples of the distribution for the Dirichlet boundary condition (D.4) determined by $\omega\sim\mathsf{Unif}([\frac{\pi}{2},7\pi])$.
@@
Expected Relative L1 $\text{Expected Relative $L^{1}$ Error}\coloneqq\operatorname{\mathbb{E}}_{a\sim\operatorname{Law}(a)}\Biggl{[}\frac{\big{\lVert}a-\mathcal{G}_{\theta}\bigl{(}(\operatorname{Id}\times\Lambda_{a})_{\#}\mu\bigr{)}\big{\rVert}_{L^{1}(\Omega)}}{\lVert a\rVert_{L^{1}(\Omega)}}\Biggr{]}\,,$ (E.9)
where $\operatorname{Law}(a)$ is the distribution of $a$ from (D.3) and $\mathcal{G}_{\theta}$ is the AMINO model. Using the same setup as in D.3, we approximate (E.9) with a Monte Carlo average over $N^{\prime}=1024$ test conductivity samples $a_{n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\operatorname{Law}(a)$ and each with $\mathsf{L}=100$ pairs of corresponding Dirichlet and Neumann boundary data. That is, the expected relative $L^{1}$ error is approximately given by the average
$\frac{1}{N^{\prime}}\sum_{n=1}^{N^{\prime}}\frac{\Big{\lVert}a_{n}-\mathcal{G}_{\theta}\Bigl{(}\tfrac{1}{\mathsf{L}}\sum_{\ell=1}^{\mathsf{L}}\delta_{(f_{\ell}^{(n)},g_{\ell}^{(a_{n})})}\Bigr{)}\Big{\rVert}_{L^{1}}}{\big{\lVert}a_{n}\big{\rVert}_{L^{1}}},\text{ where }\,\bigl{\{}(f_{\ell}^{(n)},g_{\ell}^{(a_{n})})\bigr{\}}_{\ell=1}^{\mathsf{L}}\sim\bigl{(}(\operatorname{Id}\times\Lambda_{a_{n}})_{\#}\mu\bigr{)}^{\otimes\mathsf{L}}.$ (E.10)
@@
Table 2 illustrates how OOD performance varies with the choice of training distribution. Model 1 achieves strong ID performance but performs poorly in the OOD setting. Model 2 exhibits consistent yet mediocre performance across both ID and OOD evaluations. Model 4 performs well in the ID setting but shows a noticeable drop in OOD performance. Model 3 offers the best balance, achieving low error in both ID and OOD scenarios. Notably, despite being trained on a Dirichlet data distribution with a small support regarding the parameter $\omega$, Model 3 generalizes significantly better to the larger test domain $[\frac{\pi}{2},7\pi]$ for the parameter $\omega$.
@@
In-Distribution (ID) $0.034\pm 0.019$ $0.081\pm 0.028$ $0.021\pm 0.010$ $0.017\pm 0.009$
Out-of-Distribution (OOD) $0.796\pm 0.251$ $0.078\pm 0.047$ $\mathbf{0.066}\pm 0.038$ $0.132\pm 0.047$
These results are further illustrated in Fig. 1, which presents two representative conductivity samples. The top row displays the predicted conductivities from each of the four models when evaluated on Dirichlet–Neumann data pairs sampled from the same distribution used during training. In this ID setting, all four models accurately recover the true conductivity. In contrast, the bottom row shows predictions when the Dirichlet data is drawn from the OOD setting, i.e., $\omega\sim\mathsf{Unif}([\frac{\pi}{2},7\pi])$, and the joint Dirichlet–Neumann data distribution also changes accordingly. Under this distributional shift, Models 1, 2, and 4 exhibit noticeable degradation in performance, while Model 3 maintains a reasonable level of accuracy. We use this test to demonstrate the importance of carefully selecting the training data distribution to achieve robust OOD generalization.
@@
E.3 NtD linear operator learning

### v3

@@ F.1 Bilevel minimization for function approximation
$\mathbb{Q}\coloneqq\frac{1}{K}\sum_{k=1}^{K}\delta_{\nu_{k}^{\prime}}\,.$ (F.3)
@@
The test measures $\nu_{k}^{\prime}\in\mathscr{P}_{2}(\mathbb{R}^{d})$ in (F.3) are deterministic and fixed. In applications, these could be the datasets for which we require good OOD accuracy. However, to synthetically generate the $\{\nu_{k}^{\prime}\}$ in practice, we set
@@
$\nu_{k}^{\prime}\coloneqq\mathcal{N}(m_{k}^{\prime},\mathcal{C}_{k}^{\prime})\,,\quad\text{where}\quad m_{k}^{\prime}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,I_{d})\quad\text{and}\quad\mathcal{C}^{\prime}_{k}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathsf{Wishart}_{d}(I_{d},d+1)$ (F.4)
@@
for each $k=1,\ldots,K$. Once the $K$ realizations of these distributions are generated, we view them as fixed once and for all. We assume that $K$ and $\mathbb{Q}$ are fixed, but we only have access to the empirical measure $\widehat{\mathbb{Q}}$ defined by (F.3) but with each $\nu_{k}^{\prime}$ replaced by the empirical measure $\frac{1}{M}\sum_{m=1}^{M}\delta_{v_{m}^{(k)}}$ generated by samples $\{v_{m}^{(k)}\}_{m=1}^{M}\sim{\nu_{k}^{\prime}}^{\otimes M}$. In the experiments, we set $M=5000$ and either $K=10$ or $K=1$ (a single test distribution). To create a fair test set, we use a fixed $500$ sample validation subset to update the adjoint state as in Lem. D.5 and to monitor the validation OOD accuracy during gradient descent. The remaining $4500$ samples are used to evaluate the final OOD test accuracy of our trained models; see (F.5).
Based on the available samples $\{v_{m}^{(k)}\}_{1\leq m\leq M,\,1\leq k\leq K}$ associated to the meta test distribution $\mathbb{Q}$, we construct six adaptive and nonadaptive baseline distributions for comparison:
  1. 1.
    (Normal) The Gaussian measure $\mathcal{N}(m_{0},I_{d})$ which is independent of $\mathbb{Q}$;
  2. 2.
    (Barycenter) The empirical $\mathsf{W}_{2}$ barycenter of $\mathbb{Q}$ (which is Gaussian), as described in Ex. B.8;
  3. 3.
    (Mixture) The non-Gaussian empirical mixture of $\mathbb{Q}$ given by $\frac{1}{K}\sum_{k=1}^{K}\mathcal{N}(\widehat{m}_{k},\widehat{\mathcal{C}}_{k})$, where $\widehat{m}_{k}$ and $\widehat{\mathcal{C}}_{k}$ are the empirical mean and covariance of the points $\{v_{m}^{(k)}\}_{m=1}^{M}$, respectively. Note that we enforce the mixture components to be Gaussian instead of empirical measures;
  4. 4.
    (Uniform) The $d$-dimensional uniform distribution $\mathrm{Unif}([0,1]^{d})\coloneqq\mathrm{Unif}([0,1])^{\otimes d}$, which is independent of $\mathbb{Q}$.
  5. 5.
    (nCoreSet) Nonadaptive coreset described in SM F.6 with features $\kappa({\,\cdot\,},v_{m})\in\mathcal{H}_{\kappa}$ for $v_{m}$ belonging to the fixed pool of $500K$ validation points. New indices $m$ are selected according to the maxmin sequential update (F.40) with features as above and distance between features induced by the RKHS norm of $\kappa$. We use point-by-point updates and an initial random coreset of size one.
  6. 6.
    (aCoreSet) Adaptive greedy coreset described in SM F.6 with features $\{c_{n}\kappa(u_{n},v_{m})\}_{n=1}^{N}\in\mathbb{R}^{N}$ for $v_{m}$ belonging to the fixed pool of $500K$ validation points. The numbers $\{c_{n}\}$ are the kernel method coefficients trained on points $\{u_{n}\}$, which makes this an adaptive method. New indices $m$ are selected according to the maxmin sequential update (F.40) with features as above and distance between features induced by the Euclidean norm on $\mathbb{R}^{N}$. We select new points from the pool in batches of size ten and used an initial random coreset of size six.
The pool-based active learning baselines are based on the work of Musekamp et al. [71] and results for these methods are shown in Sec. 5 and in Fig. 10.
@@
It remains to describe the optimization of the Gaussian family $\nu_{\vartheta}=\mathcal{N}(m,\mathcal{C})$ with parameter $\vartheta=(m,\mathcal{C})$. The closed form update for the mean follows from Rmk. D.9. For the covariance matrix, we optimize over the lower triangular Cholesky factor $L$ of $\mathcal{C}=LL^{\top}\in\mathbb{R}^{d\times d}$ instead of $\mathcal{C}$ itself. This removes redundant parameters due to the symmetry of $\mathcal{C}$. We enforce strict positive definiteness of $\mathcal{C}$ by replacing all nonpositive diagonal entries of $L$ with a threshold value of $10^{-7}$. Other than this constraint, each entry of $L$ is optimized over the whole of $\mathbb{R}$. Recalling the log density formula (4.5) from Ex. 4.6, we use automatic differentiation to compute $\nabla_{L}\log(p_{(m,LL^{\top})})$; this factor appears in the definition of the approximate gradient (D.11) and in Alg. 3.
@@
$\mathsf{Err}\coloneqq\sqrt{\frac{\operatorname{\mathbb{E}}_{\widehat{\nu^{\prime}}\sim\widehat{\mathbb{Q}}}\operatorname{\mathbb{E}}_{u^{\prime}\sim\widehat{\nu^{\prime}}}\big\lvert\mathcal{G}^{\star}(u^{\prime})-\widehat{\mathcal{G}}_{N}^{(\nu,\sigma)}(u^{\prime})\big\rvert^{2}}{\operatorname{\mathbb{E}}_{\widehat{\nu^{\prime}}\sim\widehat{\mathbb{Q}}}\operatorname{\mathbb{E}}_{u^{\prime}\sim\widehat{\nu^{\prime}}}\big\lvert\mathcal{G}^{\star}(u^{\prime})\big\rvert^{2}}}\approx\sqrt{\frac{\mathscr{E}_{\mathbb{Q}}(\widehat{\mathcal{G}}_{N}^{(\nu,\sigma)})}{\mathscr{E}_{\mathbb{Q}}(0)}}\,.$ (F.5)
@@
In the preceding display, we observe that $\mathsf{Err}$ employs a Monte Carlo discretization of the integrals appearing in $\mathscr{E}_{\mathbb{Q}}$ by averaging over the empirical meta measure $\widehat{\mathbb{Q}}$, which is composed of $K$ test measures. Each individual test measure is itself an empirical measure with $4500$ atoms at points that are unseen during the adjoint state calculation. See the preceding paragraph for details about our choice of $\mathbb{Q}$. We report values of $\mathsf{Err}$ in Figs. 2, 8, 10, 12, 14 and in Tab. 1.
Supplementary numerical results are given in Figs. 815. Fig. 8 reports the same information that Fig. 2 does, except now for the remaining functions $\{g_{i}\}$. The conclusions from these numerical results are mostly the same. The left column shows the training history over $50$ iterations; the “seen” curve computes the OOD error with the $\mathbb{Q}$ samples used in the bilevel algorithm, while the “unseen” curve computes the OOD error on a held-out set of $\mathbb{Q}$ samples. The third row shows a slight “generalization gap” for $\mathcal{G}^{\star}=g_{2}$ ($d=8$). The optimized Gaussian outperforms alternative distributions except for the $g_{3}$ ($d=5$) and $g_{4}$ cases, where the empirical $\mathbb{Q}$ mixture eventually overtakes it. Fig. 10 shows that this behavior is largely mitigated if $1000$ iterations of gradient descent are used to find the optimal Gaussian instead of $50$. We remark that in Fig. 2, the leftmost subplot showing the training history is designed to highlight the long-time behavior (i.e., $1000$ iterations) of the bilevel algorithm. In practice, we only run the algorithm for $O(10)$ to $O(100)$ gradient descent iterations; this is reflected in the leftmost column of Figs. 8 and 12.
While the optimized Gaussian is shown to consistently outperform nonadaptive distributions when $\mathbb{Q}$ in (F.3) is composed of $K=10$ atoms (Figs. 8 and 10) and $N$ is large enough, this is less true for a single test distribution ($K=1$), as seen in Figs. 12 and 14 when $N$ is sufficiently large. In this case, the barycenter and mixture distributions are the same test distribution and lead to trained models with lower OOD error. When $N$ is small and many iterations are used to find the optimal Gaussian, the optimal Gaussian can achieve smaller OOD error than models trained on the test distribution. Understanding for what $\mathcal{G}^{\star}$, $\mathbb{Q}$, $N$, and gradient descent iteration count makes this desirable improvement possible is an interesting and important question for future work.
To understand what properties the optimal Gaussian absorbs from $\mathcal{G}^{\star}$ and $\mathbb{Q}$, Figs. 91113, and 15 visualize the learned covariance matrices, barycenter covariances, and mixture covariances. For instance, the learned covariance puts high weight on the $x_{3}$ coordinate in the $g_{2}$ case (Fig. 11, rows 2–3, diagonal entry of matrix) even though the barycenter and mixture do not. This makes sense because $g_{2}$ is quadratic in $x_{3}$ and lower order in all other variables (recall (F.1)); thus, the learned covariance picks out the highest order term in the ground truth map. Similarly, for $g_{3}$ the learned covariance assigns high importance to the $x_{2}$ and $x_{3}$ coordinates which are also fast growing. In all cases (except for $g_{4}$), the learned covariance differs substantially from the mixture or test distribution’s covariance. These covariance visualization figures highlight an additional benefit of our $\mathcal{G}^{\star}$-dependent estimators: they lead to interpretable training distributions tailored to the structure of the problem.
Figure 8: Bilevel Alg. 3 applied to the various ground truths $\{g_{i}\}$ with $K=10$ test atoms and $50$ gradient descent iterations. (Left column) Evolution of $\mathsf{Err}$ from (F.5) vs. iterations. (Center column) $\mathsf{Err}$ of model trained on $N$ samples from optimal $\nu_{\vartheta}$ vs. nonadaptive distributions. (Right column) Same as center, except incorporating the additional function evaluation cost incurred from Alg. 1. From top to bottom, each row represents target functions $\{g_{i}\}$ corresponding to those listed in each column of Tab. 1 (from left to right). Shading represents two standard deviations away from the mean $\mathsf{Err}$ over $10$ independent runs.
Figure 9: Bilevel Alg. 3 applied with $K=10$ test atoms and $50$ gradient descent iterations. Each column visualizes the covariance matrix of the optimized (left), empirical $\mathbb{Q}$ $\mathsf{W}_{2}$ barycenter (center), and empirical $\mathbb{Q}$ mixture distributions (right). From top to bottom, each row represents target functions $\{g_{i}\}$ corresponding to those listed in each column of Tab. 1 (from left to right).
Figure 10: Bilevel Alg. 3 applied to the various ground truths $\{g_{i}\}$ with $K=10$ test atoms and $1000$ gradient descent iterations. (Left column) Evolution of $\mathsf{Err}$ from (F.5) vs. iterations. (Right column) $\mathsf{Err}$ of model trained on $N$ samples from optimal $\nu_{\vartheta}$ vs. baseline distributions. From top to bottom, each row represents target functions $\{g_{i}\}$ corresponding to those listed in each column of Tab. 1 (from left to right). Shading represents two standard deviations away from the mean $\mathsf{Err}$ over $10$ independent runs.
Figure 11: Bilevel Alg. 3 applied with $K=10$ test atoms and $1000$ gradient descent iterations. Each column visualizes the covariance matrix of the optimized (left), empirical $\mathbb{Q}$ $\mathsf{W}_{2}$ barycenter (center), and empirical $\mathbb{Q}$ mixture distributions (right). From top to bottom, each row represents target functions $\{g_{i}\}$ corresponding to those listed in each column of Tab. 1 (from left to right).
Figure 12: Bilevel Alg. 3 applied to the various ground truths $\{g_{i}\}$ with $K=1$ test atoms and $50$ gradient descent iterations. (Left column) Evolution of $\mathsf{Err}$ from (F.5) vs. iterations. (Center column) $\mathsf{Err}$ of model trained on $N$ samples from optimal $\nu_{\vartheta}$ vs. nonadaptive distributions. (Right column) Same as center, except incorporating the additional function evaluation cost incurred from Alg. 1. From top to bottom, each row represents target functions $\{g_{i}\}$ corresponding to those listed in each column of Tab. 1 (from left to right). Shading represents two standard deviations away from the mean $\mathsf{Err}$ over $10$ independent runs.
Figure 13: Bilevel Alg. 3 applied with $K=1$ test atoms and $50$ gradient descent iterations. Each column visualizes the covariance matrix of the optimized (left) and empirical test distributions (right), respectively. From top to bottom, each row represents target functions $\{g_{i}\}$ corresponding to those listed in each column of Tab. 1 (from left to right).
Figure 14: Bilevel Alg. 3 applied to the various ground truths $\{g_{i}\}$ with $K=1$ test atoms and $1000$ gradient descent iterations. (Left column) Evolution of $\mathsf{Err}$ from (F.5) vs. iterations. (Right column) $\mathsf{Err}$ of model trained on $N$ samples from optimal $\nu_{\vartheta}$ vs. nonadaptive distributions. From top to bottom, each row represents target functions $\{g_{i}\}$ corresponding to those listed in each column of Tab. 1 (from left to right). Shading represents two standard deviations away from the mean $\mathsf{Err}$ over $10$ independent runs.
Figure 15: Bilevel Alg. 3 applied with $K=1$ test atoms and $1000$ gradient descent iterations. Each column visualizes the covariance matrix of the optimized (left) and empirical test distributions (right), respectively. From top to bottom, each row represents target functions $\{g_{i}\}$ corresponding to those listed in each column of Tab. 1 (from left to right).
@@
F.2 AMINO for learning electrical impedance tomography
@@
To illustrate this, we consider the EIT problem with the AMINO solver; see SM E.3. We start with the following forward problem setup taken from Molinaro et al. [67] for EIT on the domain $\Omega\coloneqq(0,1)^{2}\subset\mathbb{R}^{2}$. The model is the two-dimensional elliptic PDE boundary value problem
$\mathopen{}\mathclose{{\left\{\begin{aligned} \nabla\cdot(a\nabla u)&=0&&\quad\text{in}\quad\Omega\,,\\ u&=f&&\quad\text{on}\quad\partial\Omega\,.\end{aligned}}}\right$. (F.6)
For one conductivity realization $a$ defined by
$x\mapsto a(x)=\exp\Biggl(\sum_{k=1}^{m}c_{k}\sin(k\pi x_{1})\sin(k\pi x_{2})\Biggr)\,$, (F.7)
where $m=\lfloor\overline{m}\rfloor$, $\overline{m}\sim\mathsf{Unif}([1,5])$, and $\{c_{k}\}\sim\mathsf{Unif}([-1,1]^{m})$, we take $\mathsf{L}\in\mathbb{N}$ different Dirichlet boundary conditions $\{f_{\ell}\}_{\ell=1}^{\mathsf{L}}$ for $f$ in (F.6). These are given by
$x\mapsto f_{\ell}(x)=\cos\Bigl(\omega\bigl(x_{1}\cos(\theta_{\ell})+x_{2}\sin(\theta_{\ell})\bigr)\Bigr)$ (F.8)
with $\omega\sim\mathsf{Unif}([\pi,3\pi])$ and $\theta_{l}\coloneqq 2\pi\ell/\mathsf{L}$ for $\ell=1,\ldots,\mathsf{L}$. We denote by $\Lambda_{a}$ the DtN operator that maps the Dirichlet data of the PDE, $f_{\ell}$, to the Neumann boundary data, $a\frac{\partial u}{\partial\mathbf{n}}|_{\partial\Omega}$, where $u$ is the PDE solution and $\mathbf{n}$ is the outward normal unit vector. Here, the Dirichlet boundary conditions are functions drawn from a probability distribution $\mu$. AMINO learns the map from the joint Dirichlet–Neumann distribution $(\operatorname{Id},\Lambda_{a})_{\#}\mu$ to the conductivity $a$.
Suppose the general distribution of interest for the Dirichlet boundary condition is generated according to the law of $f_{l}$ in (F.8) with $\omega\sim\mathsf{Unif}([\frac{\pi}{2},7\pi])$. A reasonable distribution to train upon is this very same distribution. However, suppose we were constrained to focus and train on distributions corresponding to $\operatorname{Law}(\omega)$ supported only on a subset of the domain $\frac{\pi}{2},7\pi]$. In that case, the choice of the distribution will greatly affect the OOD error when testing with $\omega\sim\mathsf{Unif}([\frac{\pi}{2},7\pi])$. To be more concrete, we will focus on four training distributions for $\omega$, which directly determines the input Dirichlet boundary condition as in SM E.3: $\mathsf{Unif}([\frac{\pi}{2},\pi])$, $\mathsf{Unif}([\frac{7\pi}{2},4\pi])$, $\mathsf{Unif}([5\pi,\frac{11\pi}{2}])$, and $\mathsf{Unif}([\frac{13\pi}{2},7\pi])$. After obtaining the well-trained neural operators (in the sense that the ID generalization error is small enough) corresponding to each training distribution, we study the OOD error of the trained neural operators for Dirichlet data obtained by drawing $\omega\sim\mathsf{Unif}([\frac{\pi}{2},7\pi])$ in (F.8).
@@
We train four distinct AMINO models; see SM E.3 to recall AMINO and SM E.1 for its DeepONet component. Each model comprises three jointly trained components:
@@
  • Trunk network: An 88–layer multilayer perceptron (MLP) with $512$ neurons per hidden layer, LeakyReLU activations, zero dropout, and a final output dimension of $p=100$.
@@
The four neural operator models “Model 1”, “Model 2”, “Model 3”, and “Model 4” are trained using four different input parameter distributions for $\operatorname{Law}(\omega)$, respectively: $\mathsf{Unif}([\frac{\pi}{2},\pi])$, $\mathsf{Unif}([\frac{7\pi}{2},4\pi])$, $\mathsf{Unif}([5\pi,\frac{11\pi}{2}])$, and $\mathsf{Unif}([\frac{13\pi}{2},7\pi])$; see (F.8) to recall how $\omega$ is reflected in the actual input distributions on function space. For each distribution, $4096$ training samples are generated. Each sample consists of $\mathsf{L}=100$ Dirichlet–Neumann boundary condition pairs with the underlying conductivity discretized on a $70\times 70$ grid for the domain $\Omega=(0,1)^{2}$. After $500$ training epochs, the trained models are evaluated on $1024$ test samples of the distribution for the Dirichlet boundary condition (F.8) determined by $\omega\sim\mathsf{Unif}([\frac{\pi}{2},7\pi])$.
@@
$\text{Expected Relative $L^{1}$ Error}\coloneqq\operatorname{\mathbb{E}}_{a\sim\operatorname{Law}(a)}\Biggl[\frac{\big\lVert a-\mathcal{G}_{\theta}\bigl((\operatorname{Id},\Lambda_{a})_{\#}\mu\bigr)\big\rVert_{L^{1}(\Omega)}}{\lVert a\rVert_{L^{1}(\Omega)}}\Biggr]\,$, (F.9)
where $\operatorname{Law}(a)$ is the distribution of $a$ from (F.7) and $\mathcal{G}_{\theta}$ is the AMINO model. Using the same setup as in E.3, we approximate (F.9) with a Monte Carlo average over $N^{\prime}=1024$ test conductivity samples $a_{n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\operatorname{Law}(a)$ and each with $\mathsf{L}=100$ pairs of corresponding Dirichlet and Neumann boundary data. That is, the expected relative $L^{1}$ error is approximately given by the average
$\frac{1}{N^{\prime}}\sum_{n=1}^{N^{\prime}}\frac{\Big\lVert a_{n}-\mathcal{G}_{\theta}\Bigl(\tfrac{1}{\mathsf{L}}\sum_{\ell=1}^{\mathsf{L}}\delta_{(f_{\ell}^{(n)},g_{\ell}^{(a_{n})})}\Bigr)\Big\rVert_{L^{1}}}{\big\lVert a_{n}\big\rVert_{L^{1}}},\text{ where }\,\bigl\{(f_{\ell}^{(n)},g_{\ell}^{(a_{n})})\bigr\}_{\ell=1}^{\mathsf{L}}\sim\bigl((\operatorname{Id},\Lambda_{a_{n}})_{\#}\mu\bigr)^{\otimes\mathsf{L}}.$ (F.10)
@@
Tab. 2 illustrates how OOD performance varies with the choice of training distribution. Model 1 achieves strong ID performance but performs poorly in the OOD setting. Model 2 exhibits consistent yet mediocre performance across both ID and OOD evaluations. Model 4 performs well in the ID setting but shows a noticeable drop in OOD performance. Model 3 offers the best balance, achieving low error in both ID and OOD scenarios. Notably, despite being trained on a Dirichlet data distribution with a small support regarding the parameter $\omega$, Model 3 generalizes significantly better to the larger test domain $\frac{\pi}{2},7\pi]$ for the parameter $\omega$.
@@
In-Distribution (ID) $034\pm 0.019 0.0$$81\pm 0.028 0.0$$21\pm 0.010 0.0$$017\pm 0.009$
[…]
These results are further illustrated in Fig. 1, which presents two representative conductivity samples. The top row displays the predicted conductivities from each of the four models when evaluated on Dirichlet–Neumann data pairs sampled from the same distribution used during training. In this ID setting, all four models accurately recover the true conductivity. In contrast, the bottom row shows predictions when the Dirichlet data is drawn from the OOD setting, i.e., $\omega\sim\mathsf{Unif}([\frac{\pi}{2},7\pi])$, and the joint Dirichlet–Neumann data distribution also changes accordingly. Under this distributional shift, Models 1, 2, and 4 exhibit noticeable degradation in performance, while Model 3 maintains a reasonable level of accuracy. We use this test to demonstrate the importance of carefully selecting the training data distribution to achieve robust OOD performance.
@@
F.3 NtD linear operator learning